OpenStudy (a1234):

Is the same reactant always the limiting reactant in a reaction, even when the masses of the reactants are changed proportionately?

OpenStudy (a1234):

@Cuanchi

OpenStudy (photon336):

well not necessarily. it all depends on how many grams of each reactant you have

OpenStudy (a1234):

I have the reaction SiO2 + 2C --> Si + 2CO. Is SiO2 always the limiting reactant in this reaction?

OpenStudy (photon336):

it depends on the mass of each reactant that you have

OpenStudy (a1234):

The sets of masses (for SiO2 and C) I'm given in grams are 380 and 820, 560 and 640, 680 and 520, and 820 and 380.

OpenStudy (ciarán95):

A balanced chemical equation tells us how much of each species (i.e. number of moles) we need to form a certain number of moles of product. So, for the reaction: \[SiO _{2} + 2C \rightarrow Si + 2CO\] we need to react 1 molecule of SiO2 with 2 carbon atoms to form 1 atom of Si and 2 atoms of CO. In other words, we need 2 moles of C for every 1 mole of SiO2 in our reaction mixture and this will form 1 mole of Si and 2 moles of CO. To find out how many moles of each reactant we actually have, we can convert their mass in grams by dividing by their relative molecular mass or relative atomic mass. \[No.~of~moles = \frac{ No.~of~grams }{ Relative~Molecular~(or~Atomic)~Mass }\]

OpenStudy (ciarán95):

If we look at the first set of figures you mentioned, where we are reacting 380g of SiO2 (Silicon Dioxide) with 820g of C (Atomic Carbon)...... A limiting reagent will be the one whose quantity limits the amount of product we can actually form - it will run out first. The other reagent will be in excess and will always have some of its starting quantity left over once the limiting reagent is all used up. How do we know if there is a limiting reagent in this case and which reactant is it? Well, we cannot directly compare 380g and 820g, but we can compare their number of moles. SiO2: Relative Molecular Mass of 23 + 2(16) = 60 g/mol \[No.~of~moles = \frac{ 380g }{ 60g/mol } = 6.33~moles\] C: Relative Atomic Mass of 12 g/mol \[No. of moles = \frac{ 820g }{ 12g/mol } = 68.3~moles \] From the balanced equation, SiO2 and C will react exactly with a molar ratio of 1:2, where there is no limiting reagent. Here, we have a molar ratio of 6.33 : 68.3, or 1 : 10.8. So, we have far more carbon present that can react with the amount of SiO2 molecules that we've added. Therefore, C is in excess which means SiO2 is the limiting reagent, limiting the amount of the products we can form. As @Photon336 said, this mightn't be the case in general though, so be careful! Hope that helps :)

likeabossssssss (likeabossssssss):

plz close this question wen u r done with it

OpenStudy (a1234):

Thank you! That was very helpful.